# Integrated Sensing and Communications in Downlink FDD MIMO without CSI Feedback

### v1

@@ III Downlink channel reconstruction & Performance Metrics
III-A Reconstructing the Downlink Channel
@@
where $\mathbf{n}_{k}\in\mathbb{C}^{NS\times 1}$ follows the additive white Gaussian noise (AWGN). Our aim is to estimate the DL channel from the UL-observed signal $\mathbf{y}_{k}$. To achieve this, we first identify the set of parameters of the UL channel $\left\{{\hat{\alpha}}^{\mathrm{ul}}_{k,\ell},{\hat{\tau}}^{\mathrm{ul}}_{k,\ell},\hat{\theta}^{\mathrm{ul}}_{k,\ell}\right\}_{\ell=1,\dots,L_{k}}$ from $\mathbf{y}_{k}$, then reconstruct the DL channel at the frequency difference $f$ by exploiting their frequency invariance discussed in Section II.
@@ III-B Communication & Sensing Performance Metric
In this subsection, we aim to characterize the communication performance metric given the imperfect CSIT, and also to define the sensing performance metric for the design of the beam pattern.
@@ III-B1 Communication Performance Metric
where $z_{k}\sim\mathcal{CN}(0,\sigma^{2})$ follows AWGN. With this signal model, we aim to maximize the ergodic sum SE under imperfect CSIT. Given that the channel estimate $\hat{\mathbf{h}}_{k}$ is provided, the instantaneous SE for the common and private messages is defined as the expectation over the CSIT error $\mathbf{e}_{k}$:
$R_{\mathrm{c},k}^{\mathrm{i}}(\mathbf{P})$ $=\mathbb{E}_{\mathbf{e}_{k}}\left[\left.\log_{2}\left(\frac{|\mathbf{h}_{k}^{\sf H}\mathbf{p}_{\mathrm{c}}|^{2}}{\sum_{i=1}^{K}|\mathbf{h}^{\sf H}_{k}\mathbf{p}_{\mathrm{p},i}|^{2}+\sigma^{2}/P}\right)\right|\hat{\mathbf{h}}_{k}\right],$ (14)
$R_{\mathrm{p},k}^{\mathrm{i}}(\mathbf{P})$ $=\mathbb{E}_{\mathbf{e}_{k}}\left[\left.\log_{2}\left(\frac{|\mathbf{h}_{k}^{\sf H}\mathbf{p}_{\mathrm{p},k}|^{2}}{\sum_{i=1,i\neq k}^{K}|\mathbf{h}^{\sf H}_{k}\mathbf{p}_{\mathrm{p},i}|^{2}+\sigma^{2}/P}\right)\right|\hat{\mathbf{h}}_{k}\right].$ (15)
@@
$R_{\mathrm{c},k}^{\mathrm{i}}(\mathbf{P})\geq\mathbb{E}_{\mathbf{e}_{k}}\left[\log_{2}(1+S_{\mathrm{c},k}I^{-1}_{\mathrm{c},k})\right]$ (19)
$\mathop{\geq}^{(a)}\log_{2}\left({1+\frac{|\hat{\mathbf{h}}_{k}^{\sf H}{\mathbf{p}}_{\mathrm{c}}|^{2}}{\begin{Bmatrix}{\sum_{i=1}^{K}|\hat{\mathbf{h}}_{k}^{\sf H}{\mathbf{p}}_{\mathrm{p},i}|^{2}+\sum_{i=1}^{K}{\mathbf{p}}_{\mathrm{p},i}^{\sf H}\mathbb{E}\left[{{\mathbf{e}}_{k}{\mathbf{e}}_{k}^{\sf H}}\right]{\mathbf{p}}_{\mathrm{p},i}}\\ +\sum_{m=1}^{M}|\hat{\mathbf{h}}_{k}^{\sf H}{\mathbf{p}}_{\mathrm{r},m}|^{2}{+{\sum_{m=1}^{M}{\mathbf{p}}_{\mathrm{r},m}^{\sf H}\mathbb{E}\left[{{\mathbf{e}}_{k}{\mathbf{e}}_{k}^{\sf H}}\right]{\mathbf{p}}_{\mathrm{r},m}}}\\ +\frac{\sigma^{2}}{P}\end{Bmatrix}}}\right)$ (20)
$\mathop{=}^{(b)}\log_{2}\left({1+\frac{|\hat{\mathbf{h}}_{k}^{\sf H}{\mathbf{p}}_{\mathrm{c}}|^{2}}{\begin{Bmatrix}{\sum_{p=1}^{K}|\hat{\mathbf{h}}_{k}^{\sf H}{\mathbf{p}}_{\mathrm{p},i}|^{2}+{\sum_{m=1}^{M}|\hat{\mathbf{h}}_{k}^{\sf H}{\mathbf{p}}_{\mathrm{r},m}|^{2}}}\\ +{\sum_{p=1}^{K}{\mathbf{p}}_{\mathrm{p},i}^{\sf H}{\bm{\Sigma}}_{k}{\mathbf{p}}_{\mathrm{p},i}+\sum_{m=1}^{M}{\mathbf{p}}_{\mathrm{r},m}^{\sf H}{\bm{\Sigma}}_{k}{\mathbf{p}}_{\mathrm{r},m}+\frac{\sigma^{2}}{P}}\end{Bmatrix}}}\right)$ (21)
@@
$R_{\mathrm{c}}=\min_{k\in\mathcal{K}}\left\{\mathbb{E}_{\hat{\mathbf{h}}_{k}}\left[R_{\mathrm{c},k}^{\mathrm{i}}(\mathbf{P})\right]\right\}$ $\geq\min_{k\in\mathcal{K}}\left\{\mathbb{E}_{\hat{\mathbf{h}}_{k}}\left[\bar{R}_{\mathrm{c},k}^{\mathrm{i}}(\mathbf{P})\right]\right\}$
$\geq\mathbb{E}_{\hat{\mathbf{h}}_{k}}\left[\bar{R}_{\mathrm{c}}^{\mathrm{i}}(\mathbf{P})\right\}.$ (23)
@@
$R_{k}^{\mathrm{i}}(\mathbf{P})\geq\mathbb{E}_{\mathbf{e}_{k}}\left[\log_{2}(1+S_{k}{I^{-1}_{k}})\right],$ (24)
$\mathop{\geq}\log_{2}\!\left({\!1\!+\!\frac{|\hat{\mathbf{h}}_{k}^{\sf H}{\mathbf{p}}_{\mathrm{p},k}|^{2}}{\begin{Bmatrix}{\sum_{i=1,i\neq k}^{K}|\hat{\mathbf{h}}_{k}^{\sf H}{\mathbf{p}}_{\mathrm{p},i}|^{2}}+\sum_{i=1}^{K}{\mathbf{p}}_{\mathrm{p},i}^{\sf H}{\bm{\Sigma}}_{k}{\mathbf{p}}_{\mathrm{p},i}\\ {+{\sum_{m=1}^{M}|\hat{\mathbf{h}}_{k}^{\sf H}{\mathbf{p}}_{\mathrm{r},m}|^{2}}+\sum_{m=1}^{M}{\mathbf{p}}_{\mathrm{r},m}^{\sf H}{\bm{\Sigma}}_{k}{\mathbf{p}}_{\mathrm{r},m}+\frac{\sigma^{2}}{P}}\end{Bmatrix}}\!}\right)$ (25)
@@
$R_{k}=\mathbb{E}_{\hat{\mathbf{h}}_{k}}\left[R_{k}^{\mathrm{i}}(\mathbf{P})\right]\geq\mathbb{E}_{\hat{\mathbf{h}}_{k}}\left[\bar{R}_{k}^{\mathrm{i}}(\mathbf{P})\right].$ (27)
@@
${R}_{\sum}\geq\bar{R}_{\sum}=\mathbb{E}_{\hat{\mathbf{h}}_{k}}\left[\bar{R}^{\mathrm{i}}_{\mathrm{c}}(\mathbf{P})+\sum_{k=1}^{K}\bar{R}^{\mathrm{i}}_{k}(\mathbf{P})\right].$ (28)
@@ III-B2 Sensing Performance Metric
As a sensing performance metric, we consider the MSE between the target beam pattern and the actual beam pattern. It should be noted that the MSE metric offers versatility, enabling features such as adaptive sensing beam widening to address angle uncertainties [30, 2, 3]. Comparisons of MSE and other possible sensing metrics such as CRLB [31] and SCNR (signal to clutter and noise ratio) are presented in [2].
@@ III-D1 Problem Reformulation
Note that the power constraint $\|{\mathbf{P}}\|_{\sf F}^{2}=1$ in $\mathscr{P}_{1}$ is also removed in $\mathscr{P}_{2}$, as this constraint will be incorporated into our proposed algorithm. Notwithstanding this reformulation, directly tackling $\mathscr{P}_{2}$ is still infeasible because evaluating SINR expressions with (III-D1)-(43) requires the error covariance matrix $\mathbf{\Sigma}_{k}$, which has not been determined. In the next section, we explain challenges in obtaining $\mathbf{\Sigma}_{k}$ and provide our approach to approximate $\mathbf{\Sigma}_{k}$.
@@ IV Error Covariance Matrix Estimation
Incorporating the error covariance matrix $\mathbf{\Sigma}_{k}$ in $\mathscr{P}_{2}$ is important to reflect knowledge of CSIT imperfections into the precoder design. In conventional approaches, we can obtain $\mathbf{\Sigma}_{k}$ through (i) Bayesian estimation or (ii) Frequentist approach. In Bayesian estimation, we assume that the channel statistics are known in the BS. For example, by modeling the DL channel as a Gaussian distribution with a known covariance, a closed-form expression for the error covariance can be derived, provided that certain CSI feedback is available [8, 25]. In the Frequentist approach, we use DL channel error samples to empirically calculate $\mathbf{\Sigma}_{k}$ as shown in [32].
@@
where $\sigma_{k,n}$ denotes the MSE at antenna $n$ for user $k$. We note that the diagonal error covariance matrix is the case in which uncorrelated channel estimation errors are encountered for each antenna, which is widely assumed for the rich scattering channel where the spatial correlation of the channel is limited [32]. By replacing $\mathbf{\Sigma}_{k}$ in (III-D1)-(43) with $\hat{\mathbf{\Sigma}}_{k}$, we are ready to solve problem $\mathscr{P}_{2}$.
@@ V Precoder Optimization
  • (i) Primal feasibility: To be a feasible solution of given problem, the precoder must satisfy the given constraint, i.e., the inequality constraint in (53).
@@
  • (iii) Complementary slackness: It refers to a relationship between the inequality constraints and their corresponding Lagrange multipliers, which represents
@@
Recall that, so far, we have assumed the given Lagrange multiplier $\nu=\nu^{\dagger}$, and our rationale holds as long as the KKT condition is satisfied. Now, to ensure this, we explore how to update $\nu$ accordingly. First, $\nu$ must be non-negative due to dual feasibility, and it must also satisfy complementary slackness when multiplied by the inequality constraints in the dual problem. Meanwhile, the gradient component related to the MSE constraint in (1), (68) is linearly proportional to $\nu$, which implies that the design objectives can be adjusted based on the value of $\nu$. We summarize this in the following remark.

### v3

@@ III Downlink channel reconstruction & Performance Metrics
III-A Reconstructing downlink channel
@@
where $\mathbf{n}_{k}\in\mathbb{C}^{NS\times 1}$ follows additive white Gaussian noise (AWGN). Our aim is to estimate the DL channel from the UL-observed signal $\mathbf{y}_{k}$. To achieve this, we first identify the set of parameters of the UL channel $\left\{{\hat{\alpha}}^{\mathrm{ul}}_{k,\ell},{\hat{\tau}}^{\mathrm{ul}}_{k,\ell},\hat{\theta}^{\mathrm{ul}}_{k,\ell}\right\}_{\ell=1,\dots,L_{k}}$ from $\mathbf{y}_{k}$, then reconstruct the DL channel at the frequency difference $f$ by exploiting their frequency invariance discussed in Section II.
@@
Remark 1.
(On the assumption of partial reciprocity): The assumption of reciprocal delay, angular, and path parameters between UL and DL channels (i.e., partial reciprocity) is valid primarily under moderate frequency separations (e.g., several hundred MHz to low GHz), as supported by prior works [21, 19]. Especially, we consider a small UL-DL frequency separation ($f\approx 190\;{\text{MHz}}\ll f_{c}^{\mathrm{ul}},f_{c}^{\mathrm{dl}}$), which results in minimal angular spread differences (less than $0.5$ degrees), as presented in 3GPP TR 38.901. For significantly larger separations (e.g., multiple GHz), the CSI reconstruction accuracy may degrade due to increased angular parameter mismatch [30]. In such cases, we can employ a Gaussian approximation technique to account for angular mismatches, enabling analytical tractability and robust performance evaluation (i.e., the generalized mutual information framework [31]).
@@ III-B Communication & sensing performance metric
In this subsection, we aim to characterize the communication performance metric given the imperfect CSI, and also to define the sensing performance metric for the design of the beam pattern.
@@ III-B1 Communication performance metric
where $z_{k}\sim\mathcal{CN}(0,\sigma^{2})$ follows AWGN. With this signal model, we aim to maximize the ergodic sum SE under imperfect CSI. Given that the channel estimate $\hat{\mathbf{h}}_{k}$ is provided, the instantaneous SE for the common and private messages is defined as the expectation over the CSI error $\mathbf{e}_{k}$:
$R_{\mathrm{c},k}^{\mathrm{i}}(\mathbf{P})$
$=\mathbb{E}_{\{\mathbf{e}_{k}\}}\left[\left.\log_{2}\left(1+\frac{|\mathbf{h}_{k}^{\sf H}\mathbf{p}_{\mathrm{c}}|^{2}}{\begin{Bmatrix}\sum_{i=1}^{K}|\mathbf{h}^{\sf H}_{k}\mathbf{p}_{\mathrm{p},i}|^{2}+\\ \sum_{m=1}^{K}|\mathbf{h}^{\sf H}_{k}\mathbf{p}_{\mathrm{r},m}|^{2}+\sigma^{2}/P\end{Bmatrix}}\right)\right|\hat{\mathbf{h}}_{k}\right],$ (14)
$R_{\mathrm{p},k}^{\mathrm{i}}(\mathbf{P})$
$=\mathbb{E}_{\{\mathbf{e}_{k}\}}\left[\left.\log_{2}\left(1+\frac{|\mathbf{h}_{k}^{\sf H}\mathbf{p}_{\mathrm{p},k}|^{2}}{\begin{Bmatrix}\sum_{i=1,i\neq k}^{K}|\mathbf{h}^{\sf H}_{k}\mathbf{p}_{\mathrm{p},i}|^{2}+\\ \sum_{m=1}^{K}|\mathbf{h}^{\sf H}_{k}\mathbf{p}_{\mathrm{r},m}|^{2}+\sigma^{2}/P\end{Bmatrix}}\right)\right|\hat{\mathbf{h}}_{k}\right].$ (15)
We confirm that the common symbol is not included in the interference term of (15) as it has already been decoded by the RSMA process. This RSMA decoding order (decoding the common message first, followed by the private message) is predetermined, which is a key feature to simplify decoder structure while achieving interference mitigation gains. For more details, we refer to [28].
@@
$R_{\mathrm{c},k}^{\mathrm{i}}(\mathbf{P})\geq\mathbb{E}_{\{\mathbf{e}_{k}\}}\left[\log_{2}(1+S_{\mathrm{c},k}I^{-1}_{\mathrm{c},k})\right]$ (19)
${\color[rgb]{0,0,0}\mathop{\geq}^{(a)}\log_{2}\left({1+\frac{|\hat{\mathbf{h}}_{k}^{\sf H}{\mathbf{p}}_{\mathrm{c}}|^{2}}{\begin{Bmatrix}{\sum_{i=1}^{K}|\hat{\mathbf{h}}_{k}^{\sf H}{\mathbf{p}}_{\mathrm{p},i}|^{2}+\sum_{m=1}^{M}|\hat{\mathbf{h}}_{k}^{\sf H}{\mathbf{p}}_{\mathrm{r},m}|^{2}}\\ {+{\mathbf{p}}_{\mathrm{c}}^{\sf H}\mathbb{E}\left[{{\mathbf{e}}_{k}{\mathbf{e}}_{k}^{\sf H}}\right]{\mathbf{p}}_{\mathrm{c}}+\sum_{i=1}^{K}{\mathbf{p}}_{\mathrm{p},i}^{\sf H}\mathbb{E}\left[{{\mathbf{e}}_{k}{\mathbf{e}}_{k}^{\sf H}}\right]{\mathbf{p}}_{\mathrm{p},i}}\\ +{\sum_{m=1}^{M}{\mathbf{p}}_{\mathrm{r},m}^{\sf H}\mathbb{E}\left[{{\mathbf{e}}_{k}{\mathbf{e}}_{k}^{\sf H}}\right]{\mathbf{p}}_{\mathrm{r},m}}+{\sigma^{2}}/{P}\end{Bmatrix}}}\right)}$ (20)
${\color[rgb]{0,0,0}\mathop{\geq}^{(b)}\log_{2}\left({1+\frac{|\hat{\mathbf{h}}_{k}^{\sf H}{\mathbf{p}}_{\mathrm{c}}|^{2}}{\begin{Bmatrix}{\sum_{i=1}^{K}|\hat{\mathbf{h}}_{k}^{\sf H}{\mathbf{p}}_{\mathrm{p},i}|^{2}+\sum_{m=1}^{M}|\hat{\mathbf{h}}_{k}^{\sf H}{\mathbf{p}}_{\mathrm{r},m}|^{2}}\\ {+{\mathbf{p}}_{\mathrm{c}}^{\sf H}{\bm{\Sigma}}_{k}{\mathbf{p}}_{\mathrm{c}}+\sum_{i=1}^{K}{\mathbf{p}}_{\mathrm{p},i}^{\sf H}{\bm{\Sigma}}_{k}{\mathbf{p}}_{\mathrm{p},i}}\\ +{\sum_{m=1}^{M}{\mathbf{p}}_{\mathrm{r},m}^{\sf H}{\bm{\Sigma}}_{k}{\mathbf{p}}_{\mathrm{r},m}}+{\sigma^{2}}/{P}\end{Bmatrix}}}\right)}$ (21)
@@
$R_{\mathrm{c}}=\min_{k\in\mathcal{K}}\left\{\mathbb{E}_{\{\hat{\mathbf{h}}_{k}\}}\left[R_{\mathrm{c},k}^{\mathrm{i}}(\mathbf{P})\right]\right\}$ $\geq\min_{k\in\mathcal{K}}\left\{\mathbb{E}_{\{\hat{\mathbf{h}}_{k}\}}\left[\bar{R}_{\mathrm{c},k}^{\mathrm{i}}(\mathbf{P})\right]\right\}$
$\geq\mathbb{E}_{\{\hat{\mathbf{h}}_{k}\}}\left[\bar{R}_{\mathrm{c}}^{\mathrm{i}}(\mathbf{P})\right\}.$ (23)
@@
$R_{k}^{\mathrm{i}}(\mathbf{P})\geq\mathbb{E}_{\{\mathbf{e}_{k}\}}\left[\log_{2}(1+S_{k}{I^{-1}_{k}})\right],$ (24)
${\color[rgb]{0,0,0}\mathop{\geq}\log_{2}\!\left({\!1\!+\!\frac{|\hat{\mathbf{h}}_{k}^{\sf H}{\mathbf{p}}_{\mathrm{p},k}|^{2}}{\begin{Bmatrix}{\sum_{i=1,i\neq k}^{K}|\hat{\mathbf{h}}_{k}^{\sf H}{\mathbf{p}}_{\mathrm{p},i}|^{2}}+{\sum_{m=1}^{M}|\hat{\mathbf{h}}_{k}^{\sf H}{\mathbf{p}}_{\mathrm{r},m}|^{2}+{\mathbf{p}}_{\mathrm{c}}^{\sf H}{\bm{\Sigma}}_{k}{\mathbf{p}}_{\mathrm{c}}}\\ {+\sum_{i=1}^{K}{\mathbf{p}}_{\mathrm{p},i}^{\sf H}{\bm{\Sigma}}_{k}{\mathbf{p}}_{\mathrm{p},i}+\sum_{m=1}^{M}{\mathbf{p}}_{\mathrm{r},m}^{\sf H}{\bm{\Sigma}}_{k}{\mathbf{p}}_{\mathrm{r},m}+\frac{\sigma^{2}}{P}}\end{Bmatrix}}\!}\right)}$ (25)
@@
$R_{k}=\mathbb{E}_{\{\hat{\mathbf{h}}_{k}\}}\left[R_{k}^{\mathrm{i}}(\mathbf{P})\right]\geq\mathbb{E}_{\{\hat{\mathbf{h}}_{k}\}}\left[\bar{R}_{k}^{\mathrm{i}}(\mathbf{P})\right].$ (27)
@@
${R}_{\sum}\geq\bar{R}_{\sum}=\mathbb{E}_{\{\hat{\mathbf{h}}_{k}\}}\left[\bar{R}^{\mathrm{i}}_{\mathrm{c}}(\mathbf{P})+\sum_{k=1}^{K}\bar{R}^{\mathrm{i}}_{k}(\mathbf{P})\right].$ (28)
@@ III-B2 Sensing performance metric
As a sensing performance metric, we consider the MSE between the target beam pattern and the actual beam pattern. It should be noted that the MSE metric offers versatility, enabling features such as adaptive sensing beam widening to address angle uncertainties [32, 2, 3]. Comparisons of MSE and other possible sensing metrics such as CRLB [33] and SCNR (signal-to-clutter and noise ratio) are presented in our previous work [2].
@@
Remark 2.
(On the choice of sensing metric) While radar-specific sensing metrics such as the CRLB and the ambiguity function offer theoretical insights into estimation accuracy and resolution, they may be less limited in certain aspects. Specifically, the CRLB requires precise knowledge of the target parameters or their distributions, which are typically unavailable during initial sensing [33]. In contrast, the beam-pattern MSE enables explicit control over angular power distribution without relying on target-specific information [2, 4]. Accordingly, it is thus well-suited for beam sweeping and adaptive sensing under uncertainty. For these reasons, beam-pattern MSE is adopted as the primary sensing metric in this work. Later, we also show that the proposed method can be extended to incorporate other radar performance metrics such as SCNR.
@@ III-D Problem reformulation
Note that the power constraint $\|{\mathbf{P}}\|_{\sf F}^{2}=1$ in $\mathscr{P}_{1}$ is also removed in $\mathscr{P}_{2}$, as this constraint will be incorporated into our proposed algorithm. Notwithstanding this reformulation, directly tackling $\mathscr{P}_{2}$ is still infeasible because evaluating SINR expressions with (III-D)-(43) requires the error covariance matrix $\mathbf{\Sigma}_{k}$, which has not been determined. In the next section, we explain challenges in obtaining $\mathbf{\Sigma}_{k}$ and provide our approach to approximate $\mathbf{\Sigma}_{k}$.
@@ IV Error Covariance Matrix Estimation
Incorporating the error covariance matrix $\mathbf{\Sigma}_{k}$ in $\mathscr{P}_{2}$ is important to reflect knowledge of CSI imperfections into the precoder design. In conventional approaches, we can obtain $\mathbf{\Sigma}_{k}$ through (i) Bayesian estimation or (ii) Frequentist approach. In Bayesian estimation, we assume that the channel statistics are known in the BS. For example, by modeling the DL channel as a Gaussian distribution with a known covariance, a closed-form expression for the error covariance can be derived, provided that certain CSI feedback is available [7, 25]. In the Frequentist approach, we use DL channel error samples to empirically calculate $\mathbf{\Sigma}_{k}$ as shown in [34].
@@
where $\sigma_{k,n}$ denotes the MSE at antenna $n$ for user $k$. We note that the diagonal error covariance matrix is the case in which uncorrelated channel estimation errors are encountered for each antenna, which is widely assumed for the rich scattering channel where the spatial correlation of the channel is limited [34]. By replacing $\mathbf{\Sigma}_{k}$ in (III-D)-(43) with $\hat{\mathbf{\Sigma}}_{k}$, we are ready to solve problem $\mathscr{P}_{2}$. We note that the robustness of our ECM estimation method can be found in our prior work [24]. As shown in Fig. 1 of [24], the estimated ECM closely matches the actual MSE and outperforms the CRLB.
@@ V Precoder Optimization
  • (i) Primal feasibility: The precoder must satisfy the given constraint, i.e., the inequality constraint in (53).
@@
  • (iii) Complementary slackness: Relationship between the inequality constraints and their corresponding Lagrange multipliers is given as
@@
Recall that we have assumed a fixed Lagrange multiplier $\nu=\nu^{\dagger}$, which is valid as long as the KKT condition holds. To ensure this, we now explore how to update $\nu$. First, $\nu$ must be non-negative due to dual feasibility, and it must also satisfy complementary slackness when multiplied by the inequality constraints in the dual problem. Meanwhile, the gradient component related to the MSE constraint in (1), (68) is linearly proportional to $\nu$, which implies that the design objectives can be adjusted based on the value of $\nu$. We summarize this in the following remark.
@@
Remark 4.
(Comparison to conventional codebook-based CSI acquisition) Following 5G NR, acquiring DL CSI in FDD systems typically involves the following steps: (i) the BS transmits CSI-RS; (ii) the user estimates the DL CSI and computes a CSI report based on a predefined codebook; (iii) the user sends this report back to the BS over the UL control channel; and (iv) the BS decodes the feedback to recover the DL CSI. This process introduces considerable latency and signaling overhead (approximately 9 ms in the worst case), which poses challenges for time-sensitive ISAC applications (e.g., autonomous driving). In contrast, our framework reconstructs DL CSI directly from UL reference signals, eliminating the need for DL training, user-side computation, and UL feedback. Nonetheless, we emphasize that our framework is not intended to completely replace existing codebook-based methods. Since it does not rely on direct CSI feedback, our approach inevitably involves CSI estimation errors, whereas codebook-based methods are generally more suitable for acquiring accurate DL CSI. However, by avoiding the significant overhead and latency, it provides a compelling alternative for scenarios demanding ultra-low latency and reduced signaling FDD ISAC applications such as autonomous driving and real-time sensing.
Remark 5.
(Key assumptions and potential limitations) We summarize key assumptions required for our method to operate efficiently. First, our method relies on the assumption of partial reciprocity between UL and DL channels. Specifically, reciprocal propagation delay, angular parameter, channel path number, and correlated DL-UL channel gains are necessary to reconstruct the DL CSI from UL reference signals. As explained in Remark 1, these hold under moderate frequency separations but may be violated under large separations [30]. Accordingly, the performance of our method degrades as the frequency separation between UL and DL increases. In addition, as 2D-NOMP is exploited, a sparse multipath environment is needed where the number of path is relatively smaller than the number of antennas and subcarriers. Further, sufficient computational resources are necessary to perform CSI reconstruction within the channel coherence time, along with sufficiently high UL SNR for reliable DL CSI reconstruction.